OpenStudy (anonymous):

Will is taking a trip to Forestville, which is 489 miles away. If he plans to spend 8 hours on the road, what speed in miles per hour, will he need to average? Round your answer to the nearest whole number. 63 mph 61 mph 60 mph 62 mph

OpenStudy (anonymous):

s = d/t s = 489/8 s = 61.125 mph

OpenStudy (anonymous):

But if you round it to the nearest whole number, 61 mph

OpenStudy (anonymous):

thanks for your help(: